# Code: https://github.com/icey-zhang/EfficientMFD

### v1

@@ 3.3 Loss Function
The total loss function is combined with an image fusion loss function $\mathcal{L}_{\mathrm{f}}$ and object detection loss $\mathcal{L}_{\mathrm{d}}$. $\mathcal{L}_{\mathrm{f}}$ consists of three types of losses, i.e., structure loss $\mathcal{L}_{\mathrm{SSIM}}$, pixel loss $\mathcal{L}_{\text{pixel }}$ and gradient loss $\mathcal{L}_{\text{grad }}$. For one fused image, it should preserve overall structures and maintain a similar intensity distribution from source images. To this end, the structural similarity index (SSIM) is introduced in function:
@@
$\begin{gathered}\mathcal{L}_{\text{pixel}}^{\text{o}}=\left\|I_{m}\circ\left(\boldsymbol{u}-max(\boldsymbol{\omega_{\mathbf{1}}}\boldsymbol{x},\boldsymbol{\omega_{\mathbf{2}}}\boldsymbol{y})\right)\right\|_{1},\\ \mathcal{L}_{\text{pixel}}^{\text{b}}=\left\|(1-I_{m})\circ\left(\boldsymbol{u}-mean(\boldsymbol{\omega_{\mathbf{1}}}\boldsymbol{x},\boldsymbol{\omega_{\mathbf{2}}}\boldsymbol{y})\right)\right\|_{1},\end{gathered}$ (8)
@@ 3.4 Gradient Matrix Task-Alignment
The VIF and OD tasks have distinct optimization objectives. VIF primarily emphasizes capturing the pixel-level relationship between image pairs, while OD incorporates object semantics within the broader context of diverse scenes. An inherent optimization barrier exists between these two tasks. We observe that the prevailing challenges in multi-task learning are arguably task dominance and conflicting gradients. We introduce a gradient matrix Task-Alignment (GMTA) by presenting the condition number to mitigate the undesired effects of optimization barrier in task-shared parameters $\boldsymbol{\theta}^{\star}$ which are supposed to be balanced between the VIF and OD tasks. The individual task gradients of VIF and OD task are calculated by $\boldsymbol{g}_{u}=\nabla_{{\theta}^{\star}}\mathcal{L}_{u}$ and $\boldsymbol{g}_{d}=\nabla_{{\theta}^{\star}}\mathcal{L}_{d}$ in the training optimization process. The gradient matrix can be defined as $\boldsymbol{G}=\left\{\boldsymbol{g}_{u},\boldsymbol{g}_{d}\right\}$. In multi-task optimization, a cumulative gradient $\boldsymbol{g}=\boldsymbol{G}\boldsymbol{w}$ is a linear combination of task gradients and the stability of a linear system is measured by the condition number of its matrix in numerical analysis. Hence the stability of the gradient matrix is equal to the ratio of the maximum and minimum singular values of the corresponding matrix: $\kappa(\boldsymbol{G})=\frac{\sigma_{\max}}{\sigma_{\min}}$. Learning from the Aligned-MTL [37], The condition number is optimal ($\kappa(\boldsymbol{G})=1$) if and only if the gradients are orthogonal and equal in magnitude. Thereby, we consider the feature learning constraint $\mathcal{S}(\boldsymbol{\theta}^{\star})$ can be defined as the following optimization to eliminate instability in the training process:
@@
Figure 3: Visualized results of image fusion on M3FD, TNO and RoadScene.
@@ 4 Experiments and Analysis
4.1 Implementation Details
We conduct all the experiments on a server with one GeForce RTX 3090 GPU. The pretrained DiffusionDet is used for the initialization of the object detection network. In the training phase, EfficientMFD is optimized by AdamW with a batch size of 1. We set the learning rate to $2.5\times 10^{-5}$ and the weight decay as $41e-$ 4. The default training iteration is only 15,000. To fully demonstrate the effectiveness of our method, we compare seven other state-of-the-art models, including VIF method (CDDFuse [51], DIDFuse [53], U2Fusion [43] PIAFusion [41], and SwinFusion [26]), and VIF-OD joint method (MetaFusion [50], and Tardal [18]). For more analyses about the other VIF-OD datasets and visualization results, see the Appendix.
4.2 Dataset
We conduct experiments on three widely-used visible-infrared image datasets: TNO, RoadScene, and M3FD [18]. TNO and RoadScene are just used to evaluate VIF performance. Besides, M3FD is adopted to evaluate both VIF and OD performance. RoadScene with 37 image pairs and TNO with 42 image pairs are only used for the VIF task in the testing stage. The VIF network is trained by the M3FD dataset. The M3FD dataset [18] which is divided into a training set (2,940 image pairs) and a testing set (1,260 image pairs) and the 300 image pairs in the testing set is used to validate the VIF task. The original image size in the dataset is $1,024\times 768$.
4.3 Metric
[…]
Table 1: Quantitative results of different fusion methods on TNO, RoadScene and M3FD datasets. The model training (Tr.) and test (Te.) time is counted on a NVIDIA GeForce RTX 3090. The best result is highlighted.
@@ 4.3 Metric
EfficientMFD 6.36 15.47 1.65 6.40 16.28 1.6 6.79 17.11 1.69 2h50m32s 0.014s
4.4 Results of Infrared-visible Image Fusion
Qualitative results of different fusion methods are depicted in Figure 3. All the fusion methods can fuse the main features of the infrared and visible images to some extent and we can clearly observe two remarkable advantages of our method. First, the significant characteristics of infrared images can be effectively highlighted by our method. As shown in Figure 3, our M3FD fusion image captures the person riding a motorcycle, and the TNO fusion image clearly represents the person. In comparison with other methods, our method demonstrates high contrast and recognition of the objects, providing not only a visually pleasing experience but also assistance for object detection tasks. Second, our method preserves rich details from the visible images, including color and texture. In Figure 3, our advantages are evident in the fusion images across the three datasets, such as the clear outline of the white car’s rear, information on the cabinet, and words and arrows on the signboard. While retaining a substantial amount of detail, our method maintains a high resolution without introducing blurriness. In contrast, other methods fail to simultaneously achieve these two advantages. Analyzed together with the results in Table 2, we can get our method accomplishes visual friendliness while effectively aiding downstream tasks. Sequentially, we provide quantitative results of different fusion methods in Table 1. Our EfficientMFD generally achieves the largest metric values. In specific, the largest average value on MI proves that our method transfers more considerable information from both two source images. Values of EN reveal that our results contain edged details and the highest contrast between objects and the background. The large VIF shows our fusion results have high-quality visual effects and small distortion compared with the source images. Moreover, our method achieves the fastest training time and only needs about 3 hours to finish the joint learning at one stage which means that faster iterative updates can be done on new datasets. The test time to generate a fused image ranked third.
Table 2: Quantitative results of object detection on M3FD datasets among all the image fusion method + detector (i.e. YOLOv5s). The best result is highlighted.
[…]
@@ 4.4 Results of Infrared-visible Image Fusion
EfficientMFD 60.1 69.5 81.4 52.2 47.6 72.2 91.8 63.8
4.5 Results of Infrared-visible Object Detection
Figure 4: Visualize results of object detection on M3FD.
@@ 4.5 Results of Infrared-visible Object Detection
To more effectively evaluate the fusion images and observe their impact on downstream detection tasks, we conduct tests using the baseline detector YOLOv5s on all SOTA methods on the M3FD dataset. We followed the same parameter settings, and the results are shown in Figure 4. The detection results are poor when using only single-modal image inputs, with instances of missed detections, such as the motorcycle and rider next to the car and people on the far right in the image. Almost all fusion methods reduce missed detections and improve confidence by fusing information from both modalities. Through the design of an end-to-end fusion detection synchronous optimization strategy, we obtained fusion images that are visually and detection-friendly, especially for occluded and overlapping objects, as seen in the blue ellipse with the motorcycle and the overlapping people on the far right in the image.
To further assess the quality of visible and infrared fusion images, we conduct a fair comparison between our method and the state-of-the-art (SOTA) methods on YOLOv5s. As shown in Table 2, the SOTA methods demonstrate an improvement in performance compared to single-modal detection, indicating that well-fused images can effectively assist downstream tasks. In contrast, our fusion method achieves better performance, providing effective prompts for object detection tasks at the image level, especially for the objects that need to be detected. With the integration of our method, the detection performance is further enhanced, reaching the highest accuracy for each class. This better demonstrates the advantages of our training paradigm.
Table 3: Effect study of gradient matrix task-alignment by comparing with different optimization methods on M3FD. The best result is highlighted.
[…]
VIF 6.09 14.90 1.48 / /
@@
VIF-OD (w/o GMTA) 6.12 14.70 1.39 90.05 62.60
VIF-OD (w GMTA) 6.36 15.47 1.65 91.80 63.83
4.6 Ablation Studies
Effect of Gradient Matrix Task Alignment. In Sec. 3.4, we introduce Gradient Matrix Task Alignment, which helps balance the advantages and conflicting gradients of VIF and OD tasks, making the training process more stable and eliminating optimization barriers between the two tasks. To verify its effectiveness, we compared separate optimizations for VIF and OD, as well as joint optimization. Specifically, VIF represents using only $\mathcal{L}_{u}$ to optimize the fusion network, OD represents using only $\mathcal{L}_{d}$ to optimize the object detection network, and VIF-OD represents simultaneous optimization of the fusion and detection networks using both $\mathcal{L}_{u}$ and $\mathcal{L}_{d}$ loss functions. The results, as shown in Table 3, indicate that methods incorporating GMTA optimization constraints with shared weights achieve the best results for both VIF and OD. This is because the VIF and OD tasks have distinct optimization objectives. VIF primarily emphasizes capturing the pixel-level relationship between image pairs, while OD incorporates object semantics within the broader context of diverse scenes. Therefore, simply optimizing the entire network with shared loss functions may be influenced by local optimal solutions of individual task, preventing the joint task from achieving the optimal solution. The accuracy of VIF-OD (w/o GMTA) in the table shows a slight decrease compared to separately training the detection network. In contrast, GMTA orthogonalizes the gradients of shared parameters corresponding to the two tasks, allowing the joint network to converge to an optimal point with fusion detection weights.
Table 4: Average results of ablation study of the number of branches in the Object-Region-Pixel Phylogenetic Tree on M3FD dataset. The best result is highlighted.
[…]
@@ 4.6 Ablation Studies
Study of Branches in the Object-Region-Pixel Phylogenetic Tree. We have implemented the Object-Region-Pixel Phylogenetic Tree to explore the hierarchical interactions under different granularity views and extract various features across multiple region scales, as introduced in Sec. 3.2. Here, we investigate the combination of the pixel feature mining module (0) and the region feature refinement module (1,2,3,4). The results are shown in Table 4. It can be observed that with the increase in the number of branches, the fusion network achieves higher image fusion quality and object detection performance. Region features provide the fusion network with multi-level semantic features of objects. However, when higher-level semantic object information is added, the performance of the fusion network declines. This is because the detailed information contained in the deeper layers of the backbone structure shared with the detection network decreases abruptly, which may affect the fusion network’s absorption of these features, thereby influencing pixel-level fusion tasks. This illustrates the importance of balancing the semantic information provided between the detection network and the fusion network with pixel-level information.
Figure 5: Visualization of task dominance and conflicting gradients in the joint learning of OD and VIF tasks.
4.7 Analysis of Gradient Matrix
As described in Section 3.4, the VIF and OD tasks pursue different optimization goals. VIF primarily focuses on capturing pixel-level relationships between image pairs, while OD integrates object semantics within the broader context of diverse scenes. To visualize the task dominance and conflicting gradients, we plot the gradient matrix in the training stage illustrated by Figure 5. We perform a GMTA operation every 1000 iteration loss updates. Blue represents the gradients of shared parameters computed by the object detection loss function, while yellow represents the gradients of shared parameters computed by the modal fusion loss function. During the training process, it can be observed that the gradient values of the object detection task are larger and dominant, while the gradient values of the fusion task are smaller. This may affect the learning process of the fusion task during training. Conversely, the utilization of GMTA effectively mitigates this gradient dominance and conflict, facilitating a balance of shared parameters between VIF and OD.

### v2

@@ 3.3 Loss Function
The total loss is combined with an image fusion loss function $\mathcal{L}_{\mathrm{f}}$ and object detection loss $\mathcal{L}_{\mathrm{d}}$. $\mathcal{L}_{\mathrm{f}}$ consists of three types of losses, i.e., structure loss $\mathcal{L}_{\mathrm{SSIM}}$, pixel loss $\mathcal{L}_{\text{pixel }}$ and gradient loss $\mathcal{L}_{\text{grad }}$. For one fused image, it should preserve overall structures and maintain a similar intensity distribution from source images. To this end, the structural similarity index (SSIM) is introduced in function:
@@
$\mathcal{L}_{\text{pixel}}^{\text{o}}=\left\|I_{m}\circ\left(\boldsymbol{u}-max(\boldsymbol{\omega_{\mathbf{1}}}\boldsymbol{x},\boldsymbol{\omega_{\mathbf{2}}}\boldsymbol{y})\right)\right\|_{1},\mathcal{L}_{\text{pixel}}^{\text{b}}=\left\|(1-I_{m})\circ\left(\boldsymbol{u}-mean(\boldsymbol{\omega_{\mathbf{1}}}\boldsymbol{x},\boldsymbol{\omega_{\mathbf{2}}}\boldsymbol{y})\right)\right\|_{1},$ (8)
@@ 3.4 Gradient Matrix Task-Alignment
The MF and OD tasks have distinct optimization objectives. MF primarily emphasizes capturing the pixel-level relationship between image pairs, while OD incorporates object semantics within the broader context of diverse scenes. An inherent optimization barrier exists between these two tasks. We observe that the prevailing challenges in multi-task learning are arguably task dominance and conflicting gradients. We introduce a Gradient Matrix Task-Alignment (GMTA) by presenting the condition number to mitigate the undesired effects of the optimization barrier in task-shared parameters $\boldsymbol{\theta}^{\star}$ which are supposed to be balanced between the MF and OD tasks. The individual task gradients of MF and OD task are calculated by $\boldsymbol{g}_{u}=\nabla_{{\theta}^{\star}}\mathcal{L}_{u}$ and $\boldsymbol{g}_{d}=\nabla_{{\theta}^{\star}}\mathcal{L}_{d}$ in the training optimization process. The gradient matrix can be defined as $\boldsymbol{G}=\left\{\boldsymbol{g}_{u},\boldsymbol{g}_{d}\right\}$. In multi-task optimization, a cumulative gradient $\boldsymbol{g}=\boldsymbol{G}\boldsymbol{w}$ is a linear combination of task gradients and the stability of a linear system is measured by the condition number of its matrix in numerical analysis. Hence the stability of the gradient matrix is equal to the ratio of the maximum and minimum singular values of the corresponding matrix: $\kappa(\boldsymbol{G})=\frac{\sigma_{\max}}{\sigma_{\min}}$. Learning from the Aligned-MTL [38], The condition number is optimal ($\kappa(\boldsymbol{G})=1$) if and only if the gradients are orthogonal and equal in magnitude. Thereby, we consider the feature learning constraint $\mathcal{S}(\boldsymbol{\theta}^{\star})$ can be defined as the following optimization to eliminate instability in the training process:
@@
Figure 3: Visual results of image fusion on M3FD.
@@ 4 Experiments and Analysis
4.1 Dataset and Implementation Details
We conduct experiments on four widely-used visible-infrared image datasets: TNO [50], RoadScene [25], M3FD [16] and DroneVehicle [2]. TNO and RoadScene are just used to evaluate MF performance. M3FD is adopted to evaluate both MF and OD performance. Besides, DroneVehicle is utilized for oriented OD. RoadScene with 37 image pairs, TNO with 42 image pairs and M3FD with 300 pairs are only used for the MF task in the testing stage, and the MF network is trained by the M3FD dataset which is divided into a training set (2,940 image pairs) and a testing set (1,260 image pairs). We conduct all the experiments with one GeForce RTX 3090 GPU, and the code of M3FD is based on Detectron2 [51], while the code of DroneVehicle is based on MMDetection 2.26.0 [52] and MMRotate 0.3.4 [53]. The pretrained DiffusionDet is used for the initialization of the OD network. In the training phase, E2E-MFD is optimized by AdamW with a batch size of 1. We set the learning rate to $2.5\times 10^{-5}$ and the weight decay as $41e-$ 4. The default training iteration is only 15,000. Different setting information about DroneVehicle can be found in the Appendix A.4 and the implementation of metric can be found in the Appendix A.1.
Table 1: Quantitative results of different fusion methods on TNO, RoadScene, and M3FD datasets. The model training (Tr.) and test (Te.) time is counted on an NVIDIA GeForce RTX 3090. The best result is highlighted.
@@ 4.1 Dataset and Implementation Details
E2E-MFD 6.36 15.47 1.65 6.40 16.28 1.60 6.79 17.11 1.69 2h50m32s 0.014s
4.2 Main Results
Results on Multimodal Image Fusion. Qualitative results of different fusion methods are depicted in Figure 3. All the fusion methods can fuse the main features of the infrared and visible images to some extent and we can observe two remarkable advantages of our method. First, the significant characteristics of infrared images can be effectively highlighted by our method. As shown in Figure 3, our M3FD fusion image captures the person riding a motorcycle. In comparison with other methods, our method demonstrates high contrast and recognition of the objects. Second, our method preserves rich details from the visible images, including color and texture. In Figure 3, our advantages are evident in the fusion images across the M3FD dataset, such as the clear outline of the white car’s rear and a man on a motorcycle. While retaining a substantial amount of detail, our method maintains a high resolution without introducing blurriness. In contrast, other methods fail to achieve these two advantages simultaneously. Sequentially, we provide quantitative results of different fusion methods in Table 1. Our E2E-MFD generally achieves the best metric values. Specifically, the largest average value on MI proves that our method transfers more considerable information from both source images. Values of EN reveal that our results contain edged details and the highest contrast between objects and the background. The large VIF shows our fusion results have high-quality visual effects and small distortion compared with the source images. Moreover, our method achieves the fastest training time to finish the joint learning at one stage which means that faster iterative updates can be done on new datasets. The test time to generate a fused image ranked third. For more visualization results, see Appendix A.5.
Table 2: Quantitative results of object detection on M3FD dataset among all the image fusion methods + detector (i.e. YOLOv5s [56]). means using the fusion images generated by E2E-MFD for object detection training. The best result is highlighted.
Task Method People Car Bus Motorcycle Lamp Truck $\text{mAP}_{50}$ $\text{mAP}_{50:95}$
@@ 4.2 Main Results
Ours (DiffusionDet) 58.5 67.7 79.9 50.3 46.2 70.2 90.3 62.1
Ours (E2E-MFD) 60.1 69.5 81.4 52.2 47.6 72.2 91.8 63.8
Figure 4: Visual results of object detection on M3FD.
Table 3: Quantitative results of object detection on DroneVehicle test sets among all the SOTA methods, where means using the fusion images generated by E2E-MFD for object detection training and testing. The best result is highlighted.
Modality Detectors Car Truck Freight Car Bus Van $\text{mAP}_{50}$
RGB RetinaNet-OBB [59] 67.5 28.2 13.7 62.1 19.3 38.1
Faster R-CNN-OBB [60] 67.9 38.6 26.3 67.0 23.2 44.6
Gliding Vertex [61] 75.8 46.1 33.8 68.1 38.7 52.5
YOLOv5s-OBB [56] 89.0 53.6 41.9 84.8 32.6 60.4
LSKNet-OBB [62] 89.5 70.0 51.8 89.4 56.9 71.5
IR RetinaNet-OBB [59] 79.9 32.8 28.1 67.3 16.4 44.9
Faster R-CNN-OBB [60] 88.6 42.5 35.2 77.9 28.5 54.6
Gliding Vertex [61] 89.2 59.7 43.0 78.8 43.9 62.9
YOLOv5s-OBB [56] 95.6 57.2 47.5 89.4 35.2 65.0
LSKNet-OBB [62] 90.3 73.3 57.8 89.2 53.2 72.8
RGB+IR UA-CMDet [2] 87.5 60.7 46.8 87.1 38.0 64.0
TSFADet [27] 89.2 72.0 54.2 88.1 48.8 70.4
CALNet [63] 90.3 76.2 63.0 89.1 58.5 75.4
Ours (YOLOv5s-OBB) 96.7 69.9 49.9 92.6 44.5 70.7
Ours (LSKNet-OBB) 90.3 77.0 63.5 89.5 59.0 75.9
Ours (E2E-MFD) 90.3 79.3 64.6 89.8 63.1 77.4
@@
Results on Multimodal Object Detection. To more effectively evaluate the fusion images and observe their impact on downstream detection tasks, we conduct tests using the baseline detector YOLOv5s on all SOTA methods on the M3FD dataset. We follow the same parameter settings, and the visualization results are shown in Figure 4. The detection results are poor when using only single-modal image inputs, with instances of missed detection, such as the motorcycle and rider next to the car and people on the far right in the image. Almost all fusion methods reduce missed detection and improve confidence by fusing information from both modalities. Through the design of an end-to-end fusion detection synchronous optimization strategy, we obtain fusion images that are visually and detection-friendly, especially for occluded and overlapping objects, as seen in the blue ellipse with the motorcycle and the overlapping people on the far right in the image. To further assess the quality of fusion images, we conduct a fair comparison between our method and the SOTA methods on YOLOv5s. As shown in Table 2, the SOTA methods demonstrate a performance improvement compared to single-modal detection, indicating that well-fused images can effectively assist downstream tasks. In contrast, the fusion image we generate has achieved the best performance on YOLOv5s. Additionally, the performance of fusion images on DiffusionDet is also impressive, albeit slightly lower than when optimizing fusion and detection tasks simultaneously. Thanks to the collaborative optimization of both tasks, the detection performance is further enhanced. Furthermore, even when compared to end-to-end object detection methods, our approach demonstrates significant performance improvements. This better underscores the advantages of our training paradigm and the effectiveness of our method. For further visualization results, please refer to the Appendix A.6.
Results on Multimodal Oriented Object Detection. As shown in Table 3, our fusion detection synchronous optimization strategy achieves the highest accuracy. Furthermore, the outstanding detection performance on YOLOv5s-OBB [56] and LSKNet [62] using the generated fusion images (with at least 5.7% and 3.1% higher AP values compared to single modalities) demonstrate the robustness of our method. This validates the superior quality of the fusion images, indicating that they are not only visually appealing but also provide rich information for the detection task. For further visualization results, please refer to the Appendix A.4.
4.3 Ablation Studies
Analysis of Gradient Matrix. As described in Section 3.4, the MF and OD tasks pursue different optimization goals. To visualize the task dominance and conflicting gradients, we plot the gradient matrix in the training stage illustrated by Figure 5. We perform a GMTA operation every 1,000 iteration loss updates. Blue represents the gradients of shared parameters computed by the OD loss function, while yellow represents the gradients of shared parameters computed by the MF loss function. During the training process, it can be observed that the gradient values of the OD task are larger and dominant, while that of the MF task are smaller. This may affect the learning process of the fusion task during training. Conversely, the utilization of GMTA effectively mitigates this gradient dominance and conflict, facilitating a balance of shared parameters between MF and OD.
Task EN MI VIF $\text{mAP}_{50}$ $\text{mAP}_{50:95}$
MF 6.09 14.90 1.48 / /
@@ 4.3 Ablation Studies
E2E-MFD (w/o GMTA) 6.12 14.70 1.39 90.05 62.60
E2E-MFD (w GMTA) 6.36 15.47 1.65 91.80 63.83
Table 4: The validation of GMTA on M3FD.
Effect of Gradient Matrix Task-Alignment. To verify the effectiveness of GMTA, we compare separate optimizations for MF and OD, as well as joint optimization, w/o and w indicating whether to use GMTA. Specifically, MF represents using only $\mathcal{L}_{u}$ to optimize the fusion network, OD represents using only $\mathcal{L}_{d}$ to optimize the object detection network, and E2E-MFD represents simultaneous optimization of the fusion and detection networks using both $\mathcal{L}_{u}$ and $\mathcal{L}_{d}$ loss functions. The results, as shown in Table 4, indicate that methods incorporating GMTA optimization constraints with shared weights achieve the best results for both MF and OD. This is because MF primarily emphasizes capturing the pixel-level relationship between image pairs, while OD incorporates object semantics within the broader context of diverse scenes. Therefore, optimizing the entire network with shared loss functions may be influenced by local optimal solutions of individual tasks.The accuracy of E2E-MFD (w/o GMTA) shows a slight decrease compared to separately training the detection network. In contrast, GMTA orthogonalizes the gradients of shared parameters corresponding to the two tasks, allowing the joint network to converge to an optimal point with fusion detection weights. For more comparison of different multi-task Learning, please refer to Appendix A.3.
Figure 5: Visualization of task dominance and conflicting gradients in joint learning of OD and MF.
Branch EN MI VIF $\text{mAP}_{50}$ $\text{mAP}_{50:95}$
@@
Table 5: Ablation study of the number of ORPPT branches on M3FD.
Study of Branches in the Object-Region-Pixel Phylogenetic Tree. We investigate the combination of the pixel feature mining module (0) and the region feature refinement module (1,2,3,4). The results are shown in Table 5. It can be observed that with the increase in the number of branches, the fusion network achieves higher image fusion quality and object detection performance. Region features provide the fusion network with multi-level semantic features of objects. However, when higher-level semantic object information is added, the performance of the fusion network declines. This is because the detailed information contained in the deeper layers of the backbone structure shared with the detection network decreases abruptly, which may affect the fusion network’s absorption of these features, thereby influencing pixel-level fusion tasks. The different feature map visualization can be seen in Appendix A.2.